Mr. Butts smoked 15 cigarettes per day. Whats the probability a smoker will smoke less than Mr. Butts. mean = 11. standard deviation = 5
Answer by stanbon(75887)

Mr. Butts smoked 15 cigarettes per day. Whats the probability a smoker will smoke less than Mr. Butts. mean = 11. standard deviation = 5
-------------------------------------------
z(15) = (15-11)/5 = 0.8
-----
P(x<15) = P(z<0.8) = 0.7881
==============
Cheers,
Stan H.